OpenStudy (anonymous):

In a certain clothing store, 3 shirts and 5 ties cost $60, and 2 shirts and 3 ties cost $39. What is the cost of each shirt? If s = the cost of a shirt and t = the cost of a tie, which system of equations represents the problem? 3t + 5s = 60 and 2t + 3s = 39 3s + 5t = 60 and 2s + 3t = 39 s + t = 60 and 5s + 8t = 99 B

OpenStudy (anonymous):

Create tow equations: s+t=60 s+t=39 Now place the number bought respectively with each variable: 3s+5t=60 2s+3t=39 Do you know it from here or do you want me to continue explaining?

OpenStudy (anonymous):

two*

OpenStudy (anonymous):

continue

OpenStudy (anonymous):

?

OpenStudy (anonymous):

did you awnser it or no ???

OpenStudy (anonymous):

Let me try and finish what turtleluver started, Start with 3Shirts + 5Ties = 60 and 2Shirts + 3Ties = 39 understand?

OpenStudy (anonymous):

so wahts the awsner

OpenStudy (anonymous):

Not really, what you first have to do is work out how much one of these either shirts or ties costs. you need to take the top equation and times it by 2 and the bottom equation and times it by 3 do this and what do you get??